6s PROBLEM 1. what an ount will be owed in S year, i, SS00 browed now at 10% per year simple interest Answer 57,300 rd:: For I need steps for the answer
Answer #1

Total simple interest for N years = Principal x Number of years x Interest rate

(1)

Interest for 5 years ($) = 5,000 x 5 x 10% = 2,500

Amount owed ($) = Principal + Interest = 5,000 + 2,500 = 7,500

(2)

Interest for the period ($) = 625 - 500 = 125

If number of years be N, then

125 = 500 x N x 10%

125 = 50 x N

N = 2.5 years

(3) Number of years is hazy and unreadable.

(4)

With simple interest,

Interest after 5 years ($) = 5,000 x 5 x 0.15 = 3,750

With compound interest,

Total amount after 5 years ($) = 5,000 x (1.15)5 = 5,000 x 2.0114 = 10,057

Interest ($) = Amount - Principal = 10,057 - 5,000 = 5,057

Therefore, compound interest is higher by $(5,057 - 3,750) = $1,307.

NOTE: As per Answering Policy, 1st 3 questions are answered.